OpenStudy (anonymous):

Prove that x^4+x^2=1

OpenStudy (anonymous):

whats x ?

OpenStudy (anonymous):

variable :P

OpenStudy (anonymous):

i mean is it cube root of unity ? or imaginary number i or something else ? if not, then u just said its variable, so x^4+x^2 can't equal 1.

OpenStudy (anonymous):

or do u need to solve for x ???

OpenStudy (anonymous):

I don't know that is what I am trying to figure out... this was the question in my paper and I left it blank

OpenStudy (anonymous):

just prove it to 1

OpenStudy (anonymous):

by means of anything

OpenStudy (anonymous):

good luck with that :)

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

i think question is incomplete...we can solve for x, but can't prove that =1...

OpenStudy (anonymous):

PROVE ???

OpenStudy (anonymous):

x= 1 :P

OpenStudy (anonymous):

nopes, x not =1 :P

OpenStudy (anonymous):

then x= -1 :P

OpenStudy (anonymous):

nopes, x not =-1 :P

OpenStudy (anonymous):

x^4 +x^2=1 => => x^4 +x^2 -1 = 0 => (x +0.786)*(x -0.786)*(x^2 +1.618) = 0 Real solutions: Root 1: -0.786 Root 2: 0.786 Complex roots: Root 3: 0+1.272 * i Root 4: 0-1.272 * i

OpenStudy (anonymous):

are u solving that for x ,saura ?

OpenStudy (anonymous):

there you go ^^

OpenStudy (anonymous):

u copied that ^

OpenStudy (anonymous):

yeah should i give the source ? :P

OpenStudy (anonymous):

it is equal to 1

OpenStudy (anonymous):

take a screenshot of your question

OpenStudy (anonymous):

Looks like no solution